OpenStudy (anonymous):

nc0 + nc3 + nc6 ......Please help

OpenStudy (anonymous):

it goes on until nCn

OpenStudy (anonymous):

expand (1+ x)^n bionomialy then put x = the cube roots of unity you'll get 3 equations which on adding and subtracting will allow you to create the series you want

OpenStudy (anonymous):

p.s. what courses are you studying?

OpenStudy (anonymous):

ive just graduated from year 12

OpenStudy (anonymous):

neat.....which college/university/institute are you trying for

OpenStudy (anonymous):

i think ill get into one of the iits

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

so r u indian?

OpenStudy (anonymous):

wt are you doing??

OpenStudy (anonymous):

i dont see an answer

OpenStudy (anonymous):

hes right...that druvidfae is bloody bright

OpenStudy (anonymous):

z^3 = 1 , how does that help ?

OpenStudy (anonymous):

u put in the cube roots of unity one by one and add all three equations...the required series comes out

OpenStudy (anonymous):

because all other binomial coefficients like nc1 and nc2 and so on have a (1 + w + w^2) multiplied

OpenStudy (anonymous):

get it??

OpenStudy (anonymous):

ok so ( 1 + z^1/3 ) ( 1 + z^1/3)( 1 + z ^1/3)

OpenStudy (anonymous):

but roots of unity are not real

OpenStudy (anonymous):

yes..first put in 1, then w and then w^2..u dont have to use only real nos ;)

OpenStudy (anonymous):

one sec, im a bit rusty here, ( 1 + x ) ^m = 1 + mx + m(m-1)/2 x^2 + ..

OpenStudy (anonymous):

same way if i wanted nC0 + nC5 + nC10... i would have put in the fifth roots of unity..

OpenStudy (anonymous):

sorry had to leave for a moment

OpenStudy (anonymous):

hey druvidfae wtr u studying>?

OpenStudy (anonymous):

but n is not given?

OpenStudy (anonymous):

the answer is in terms of n

OpenStudy (anonymous):

yes you would the problem...is that you would have lrage number of equations also in other situations only substitutions for a prime no. will work.

OpenStudy (anonymous):

Though I'm doing my degree in computer science I am interested in all branches of science and mathematics ... particular abstract maths(algebra,m-th,etc)

OpenStudy (anonymous):

oh wait, so youre using the fact that ( 1 + w + w^2 ) = ...

OpenStudy (anonymous):

where w^3 = 1

OpenStudy (anonymous):

= 0

OpenStudy (anonymous):

1 + n + n(n-1)/2 *1 +.... 1 + nw + n(n-1)/2 *w + ... 1 + nw^2 + (n)(n-1)/2 w^2 + ...

OpenStudy (anonymous):

woops

OpenStudy (anonymous):

Though I'm doing my degree in computer science I am interested in all branches of science and mathematics ... particular abstract maths(algebra,m-th,etc)

OpenStudy (anonymous):

since the roots repeat themselves modularly i.e.after every 3 powers. the sequence effectively has only three variable inserting subsequent values allows the offset to change thus allowing us to affect every third element symmetrically

OpenStudy (anonymous):

1 + n + n(n-1)/2 *1 +.... 1 + nw + n(n-1)/2 *w^2 + ... 1 + nw^2 + (n)(n-1)/2 w^4 + ..

OpenStudy (anonymous):

ok, and w^4 = w , that sort of thing

OpenStudy (anonymous):

whats the left side of this expression?

OpenStudy (anonymous):

do you have an equation ?

OpenStudy (anonymous):

i get 3 + 0 + 0 + ...

OpenStudy (anonymous):

its similar to interference in physics .Waves with a particular offset (phi) affect the the superposition in a certain way (standing waves, packets ,etc) here we are changing the wavelength (period) rather than the offset to match our sequence.once we've got our wavelength we continue to change the offset to get the shape we need.

OpenStudy (anonymous):

canto:yes... the principle can be extended to other problems as well.

OpenStudy (anonymous):

whoa whoa , slow down there buddy

OpenStudy (anonymous):

first thing is first, youre using the identity ( 1 + 1) ^n = Sum nCk , from k=0 to n ?

OpenStudy (anonymous):

you guys leave out too many steps.

OpenStudy (anonymous):

youre doing ( 1 + w)^n, (1+1)^n, and ( 1 + w^2) ^n, adding them

OpenStudy (anonymous):

for this question we don't need the 2^n identity

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

( 1 + w)^n + (1+1)^n, + ( 1 + w^2) ^n = 3 + n( 1 + w + w^2) + nC2 ( 1 + w^2 + w^4 ) + ...

OpenStudy (anonymous):

( 1 + w)^n + (1+1)^n + ( 1 + w^2) ^n = 3 + n( 1 + w + w^2) + nC2 ( 1 + w^2 + w^4 ) + nC3 ( 1 + w^3 + w^8 ) + ... nCn ( 1 + w^n + w^(2^n))

OpenStudy (anonymous):

IS THAT WHAT YOU ARE DOING ????

OpenStudy (anonymous):

and then stuff vanishes on the right side, but the left side i dont know what it is

OpenStudy (anonymous):

the left side is a closed formula?

OpenStudy (anonymous):

i would like you guys to actually solve it, lol

OpenStudy (anonymous):

im still learning this stuff, sorry

OpenStudy (anonymous):

i messed up

OpenStudy (anonymous):

( 1 + w)^n + (1+1)^n + ( 1 + w^2) ^n = 3 + n( 1 + w + w^2) + nC2 ( 1 + w^2 + w^4 ) + nC3 ( 1 + w^3 + w^6 ) + ... nCn ( 1 + w^n + w^(2n))

OpenStudy (anonymous):

yes...as for the left side notice that the w and w^2 factors have imaginary parts which are inverse of one another expanding this using euler's form will reveal that the imaginary parts get canceled out.

OpenStudy (anonymous):

ok , im a little worried about the right side though

OpenStudy (anonymous):

= 3 + 0 + 0 + nC3 + 0 + 0

OpenStudy (anonymous):

why all non-3 get reduced to 0 . that is the sequence we want

OpenStudy (anonymous):

( 1 + w)^n + (1+1)^n + ( 1 + w^2) ^n = 3 + 3*n( 1 + w + w^2) + 3*nC2 ( 1 + w^2 + w^4 ) + 3*nC3 ( 1 + w^3 + w^6 ) + ... 3*nCn ( 1 + w^n + w^(2n))

OpenStudy (anonymous):

so the final answer is [2^n + (1+w)^n + (1+w^2)^n ]/3 = nC0 + n C 3 + ... n C n

OpenStudy (anonymous):

although n is not necessarily a multiple of 3

OpenStudy (anonymous):

why all non-3 get reduced to 0 . that is the sequence we want

OpenStudy (anonymous):

im just checking , do you agree with my closed form formula ?

OpenStudy (anonymous):

also we have to assume that n | 3

OpenStudy (anonymous):

errr, 3 | n

OpenStudy (anonymous):

yeah the l.h.s. reduces to something like 3^n (note there are probably some more factors i'm missing)

OpenStudy (anonymous):

hmmm, no that doesnt make sense, it should be less than 2^n

OpenStudy (anonymous):

since nC0 + nC1 + ... n Cn = 2^n

OpenStudy (anonymous):

oh you mean before dividing by 3 ?

OpenStudy (anonymous):

if n is not a multiple of the then the last 1 or 2 factors do not appear in the final series........since series is up to n.

OpenStudy (anonymous):

we have deletion of terms

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

then you would zeroes as the last 1 or two terms ?

OpenStudy (anonymous):

ok im going to test this , offhand i dont know the third root of unity of 1

OpenStudy (anonymous):

e^ ( i * pi/3 ) ?

OpenStudy (anonymous):

no... \[w =(1+ i \sqrt{3})/2 ; w^2 = (1-i \sqrt{3})/2\] expand both bionomially notice that odd powers cancel and the remaining terms are of the form 3^n

OpenStudy (anonymous):

1 = e^(2*Pi *i) , so 1^(1/3) = e^(2pi/3 *i)

OpenStudy (anonymous):

yes your forms are also correct

OpenStudy (anonymous):

sorry, i have to start from fundamental principles, i have horrible memory

OpenStudy (anonymous):

1 = e^( i*[ 2*Pi + 2pi *n ) , so 1^(1/3) = e^(i * [2pi/3 +2pi/3*n)

OpenStudy (anonymous):

so 1^1/3 = e ^ ( i * (2pi/3 , 4pi/3

OpenStudy (anonymous):

oh its easier if we start with theta = 0

OpenStudy (anonymous):

1 = e^( i*[ 0+ 2pi *n ) , so 1^(1/3) = e^(i * [2pi/3*n)

OpenStudy (anonymous):

so we have z^(1/3) = { 1 , e^(i*2pi/3) , e^ ( i*4pi/3)

OpenStudy (anonymous):

i assume you have it memorized down well :) so i agree with that

OpenStudy (anonymous):

so w and w^2 are conjugates, interesting

OpenStudy (anonymous):

but it might be more succinct calculator wise to use exponential polar form

OpenStudy (anonymous):

[2^n + (1+e^(i*2pi/3))^n + (1+e^(i*4pi/3))^n ]/3 = nC0 + n C 3 + ... n C n

OpenStudy (anonymous):

now test, let n = 8

OpenStudy (anonymous):

sorry my early values have a slight correction the real part should be negative.

OpenStudy (anonymous):

let n = 8 , so we have [2^8 + (1+e^(i*2pi/3))^8 + (1+e^(i*4pi/3))^8 ]/3 = 8C0 + 8 C 3 + ... 8C8 ?

OpenStudy (anonymous):

or last term should be 8 C 6

OpenStudy (anonymous):

does your calculator allow complex calculations......i.e. calculations involving iota based calculations.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get 85 for left side 85 + 3.3666666666666666666666E -13i

OpenStudy (anonymous):

the last term should be\[(n & 6)\]